OpenStudy (anonymous):

simplify the expression 7/7-2/28

OpenStudy (anonymous):

find the lowest common denominator for both fractions, then convert both and subtract the numerators!

OpenStudy (shamim):

7/7-2/28 =(28-2)/28 =26/28=13/14